EVA Report #10 – December 2nd

Crew 216 EVA Report 02-Dec-2019
EVA # 10
Author: Whittle

Purpose of EVA: Climb to summit of North Ridge to determine accessible route and identify site for potential repeater site.

Start time: 0830
End time: 1000

Narrative: Crew left Hab on foot. Arrived at North Ridge. Climbed high point on North Ridge at grid 12 S 0518464 4251538. The route taken was via the South West ridge. Time of day was ideal as ground was frozen (and only slightly thawed on decent). Would potentially be hazardous later in the day. Excellent rebroadcast location on top of the ridge. Marked on GPS, along with track of route up (available on request). Fogging was an issue on climb. In visual contact with Hab at all times. Returned to Hab via same route.

Destination: North Ridge
Coordinates (use UTM NAD27 CONUS): 12 S 0518464 4251538.

Participants: Whittle, Kennedy, Levesque

Road(s) and routes per MDRS Map:
1) Walk to North Ridge from Hab
2) Ascend highpoint at 12 S 0518464 4251538 via South West ridge
3) Return to Hab via same route

Mode of travel: Walk

EVA Report – December 2nd

Crew 216 EVA Report 02-Dec-2019
EVA # 11
Author: Kennedy
Purpose of EVA: Gather footage for education (Ho) and documentary (Alexandrova) from dinosaur quarry site.

Start time for EVA: 1330
End time for EVA: 1615

Narrative: Crew left airlock at approx 1335 and left on rovers Spirit and Curiosity, both registering at 100% battery. Road was quite muddy in spots but very manageable. This site is the greatest distance we have yet traveled, as such we minimized excessive accelerations to maximize battery life. We arrived at the parking area for the dinosaur quarry with 62% and 63% battery on our rovers. We examined the site, the quarry has been covered in. Dirt this time of year so nothing to note. We traveled along the ridge to the overlook at 12s 0518250 and 4256590. Michael set up his drone and did aerial photos of the crew. Evgenia used her tripod and took several shots. The canyon had a very distinct shape and rock characteristic that is unique from our other sites. We were quite happy with the outcome of the EVA and feel like we got excellent footage. We returned to our vehicles at 1530 and returned to the airlock in the Hab at 1615.

Coordinates (use UTM NAD27 CONUS): 12 S 0518400 4255800.

EVA Participants: Kennedy, Alexandrova, Ho

Road(s) and routes per MDRS Map:
1) North from Hab to 12 S 0518400 4255800 via Cow Dung Road
2) Explore quarry site and collect footage
3) Return to Hab via same route

Mode of travel: Rover
Vehicles used: Curiosity, Spirit

Supplemental Operations Report – December 2nd

MDRS Supplemental Operations Report 2-Dec-2019

Name of person filing report: Shannon Rupert

Reason for Report: Routine update

Non-nominal systems: RAM electrical

Action taken for non-nominal systems: Troubleshooting, no solutions. We now have our second incidence of electrical work done without permission and with no reportage. Someone has rewired the east side electric in the RAM that had been incorrectly wired when it was built. In the meantime, this appears to have damaged the wiring to the rover chargers and now they are not functional, and continually trip when the rovers are plugged in. This appears to be an issue separate from the initial problem of incorrectly installed GFI sockets and insufficient breaker strength. We intend to take up this issue in the next few weeks.

People doing things without notifying us is a common problem. Yesterday the power system had been adjusted, but not reset, so I was able to easily readjust it back where is needed to be. Later it was discovered that the crew did not have water because someone in the crew had turned it all off. But no one will say that they did it. It is the same reluctance to speak up that happened when the heat in the GreenHab was turned off, it was physically impossible for it to have happened without someone doing it, but no one wanted to say that they had done it. Until we get to the point where everyone is willing to practice full disclosure, we cannot and will not have a safe and successful Mars mission. This is something we need to get used to doing here.

Generator check, note if oil and coolant added: coolant added. Suspect
oil leak appears to not be a problem.

ScienceDome AC unit, on and operational: nothing to report

Solar— Running nominally with nighttime generator backup

Diesel Reading – % not checked

Propane Reading – not checked

Ethanol Free Gasoline – n/a gallons. Now stored in the shed.

Water (director’s tank) – n/a gallons

Water (static tank) – n/a gallons

Water (GreenHab) – n/a gallons

Water (loft) – nothing to report

Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: nothing to report

Deimos rover used: Yes, left today for repairs in Richfield

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Yes, operation nominal

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes, operation nominal

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Yes, left today for brake repairs in Richfield

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Yes, operation nominal.

Hours:

Beginning charge:

Ending charge:

Currently charging:

Notes and action taken on rovers: Rovers were cleaned of mud by the crew.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). No. 300 still in town.

Reason for use: n/a

Oil Added? n/a

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used: none

Notes and action taken on ATVs: n/a

HabCar running? If used, why, where? Yes, used to town for water
runs and supplies

CrewCar running, If used why, where? Yes, currently onsite and used to transport crews. Noted that the check engine light and service tire monitor light were not serviced when the vehicle was in the shop. Trying to determine if we paid for that or not, will probably return to dealer in Grand Junction if not.

Campus wide inspection, if action taken, what and why? Burn barrel was finally filled with holes and will be put into service back in the Outpost so we won’t have to be in sight of the crew. Old burn site will be restored to pre-burn status as soon as possible.

Evidence of rodents, where and action taken: We appear to have one brave little beast living in Little Outpost (formerly the Intern Trailer). So far he has just mocked Atila’s attempts at capture and keeps eating the cookies.

Summary of internet: All nominal.

Summary of suits and radios: Changes to radio channels has made a great improvement in COMMS.

Summary of Hab operations: New water heater has been installed. This is not an on demand model. Those have proven to be unsuited to our water system. This is a 6 gallon RV tank model, installed on the lower deck above the rear airlock, where a water heater had been
installed before. It is both gas and electric. Testing it in electric mode tonight and tomorrow we will install the 12 V power and switch for propane.

Propane was reconnected to the stove and hot plate was removed.

Grey vents from former hot water heater were reinstalled on the exterior of the Hab, covering the two holes.

Water tank heater installed and operational.

Summary of Outpost operations: Emptying the blackwater tank of the Little Outpost the first time was only marginally successful. We need to install a permanent line. Shower drain repaired.

Septic needs to be pumped. Riser was affixed to the tank so yard could be leveled and hole filled in. It will take a while to move all the dirt back into place but we have started on it.

Tank heater operational.

Summary of GreenHab operations: 40% shade cloth covering 50 % of Greenhab south side. A 30% shade cloth was ordered to cover the other 50%.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Electrical work required. See non-functional systems report above

Summary of any observatory issues: Atila and Peter are troubleshooting a cable issue on the robotic observatory. Testing will determine if it has been resolved.

Summary of health and safety issues: We do not have a functional blood pressure monitor so I ordered one to arrive this week.

List and summary of projects:

Tennis balls to protect the tarps on the tunnels: Let’s just say we hate cutting tennis balls. We vow to finish soon.

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report – December 2nd

Crew 216 Operations Report 02-Dec-19
SOL: 8
Name of person filing report: Mike Lawson

Non-nominal systems: NA
Generator: run
Hours run: 12.25
From what time last night: 1845
To what time this morning: 0700
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 52 %

Diesel Reading – 36
Station Propane Reading – 54
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1464017
Water (static tank) (gal): 269
Static to Loft Pump used – yes
Water in Green Hab (gal): 255
Water in Science Dome (gal): 0

Toilet tank emptied: no

Deimos rover used: Off Campus for repairs
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: yes
Hours: 115.7
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 38%
Currently charging: yes

Opportunity rover used: Off Campus for repairs
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging:

Curiosity rover used: yes
Hours: 120.7
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 35.%
Currently charging: yes

Notes on rovers: Opportunity to be moved off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where?
CrewCar used and why, where?

General notes and comments: EVA North Ridge, radio relay site surveyed. EVA to vicinity quarry.

Summary of internet: No serious issues.

Summary of suits and radios: No issues.

Summary of Hab operations: Routine maintenance and housekeeping tasks. Replacement of hot water tank installed, running in 120 VAC mode only. Needs an additional switch to run on propane. Will not be an issue when the station is on the generator. Generator frequency is slightly off, electric clocks run fast by 10-15 minutes/day. Not an issue when on the battery bank.

Summary of GreenHab operations: routine watering.

Summary of Science Dome operations: NA

Summary of RAM operations: RAM Rover power charging stations inoperable, electrical issues remain unresolved. Director is aware of issue.

Summary of any observatory issues: NA

Summary of health and safety issues: NA

Questions, concerns and requests to Mission Support: 2 bag NB, 1 bag B trash in the rear airlock.

Astronomy Report – December 2nd

Crew 216 Astronomy Report 02-DEC-19
Crew Astronomer: Michael Ho

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: MDRS-WF
Objects to be Imaged this Evening: The Great Orion Nebula
Images submitted with this report: nil
Problems Encountered: nil

MUSK OBSERVATORY
Solar Features Observed: None.
Images submitted with this report: nil
Problems Encountered: Nil.

GreenHab Report – December 2nd

Crew 216 GreenHab Report 02-DEC-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 12.40 pm, closed at 5.10 pm.
Average temperature: 18C
Hours of supplemental light: Automatic system
Daily water usage for crops: 16L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 09.20 am; 4.40 pm.

SEASONAL PLANTING HISTORY:
(Plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)

Crew 216:
26/11/2019:
1 avocado seed in small black pot
24/11/2019:
1 avocado seed in small black pot
23/11/2019:
seeded Zucchini in 2 large black pots
seeded Chilli Pepper ‘Shishito’ in 2 2-gallon black pots
seeded Sweet Greenfield Pepper ‘California Wonder’ in 2 2-gallon black pots
seeded Swiss Chard ‘Flamingo Pink’ in 4 2-gallon black pots
transplanted three self-seeded tomato plants in large blue tub transplanted to 2-gallon pots

OTHER NOTES:
Vents of heater readjusted back to downward-facing to prevent lower plants from freezing.
All 3xSnow peas and 3xCucumbers were relocated to metal rack, on the same tier as the Rosemary ‘Barbecue’.
The 3xStrawberries were relocated to directly opposite from where they were.

Harvest: None
Support/supplies requested:nil.

Journalist Report – December 2nd

Crew 216 Journalist Report 02-DEC-19
Sol 8
Author: Evgenia Alexandrova

On Mars you have all the time in the world to think through the philosophical questions that you never have time to explore on Earth. Notions like happiness and meaning of life are the most perplexing ones. What comes first? Meaning or happiness? Austrian psychiatrist Viktor Frankl survived concentration camps during the WWII and wrote a book based on his experience. He states that meaning is the crucial thing that drives a human to continue living through a whatever challenge. The faith in ultimate goal, the hidden or obvious sense in life can lead you through the darkest times.

People who will colonize Mars will definitely have a profound sense of meaning. I felt a glimpse of it during the three interviews with the crew members I conducted during the past days. What about MDRS? Is there a meaning in this place? Definitely. But to me, it goes back to happiness. Happiness is meaningful just by itself.

New anonymous poem of the day:

Today we saw amazing things

And felt the wonder that it brings

To feel what it is like in space

But not forget this special place

[end[

Commander's Report – December 2nd

Crew 216 Commander Report 02-DEC-2019

Sol: 8

Summary Title: Progress on Projects

Author: Marc Levesque

Crew 216 accomplished one of its main objectives during its mission by reaching the summit of the North Ridge to confirm there is an accessible route to the summit and that it would be an ideal site for a radio repeater. This repeater would be the heart of an enhanced communication system that would allow the Hab to remain in direct radio contact with EVA teams over a much wider area around MDRS, thereby enhancing EVA safety. The second EVA of the day captured some spectacular drone footage of an EVA team walking in Lithe Canyon for Michael’s project and allowed Evgenia to film additional footage to complement what she already has for her project. In a discussion over this evening’s dinner, the only group meal that seems to range from simple to complex, we discussed the optimal size for a crew at MDRS, or on a Mars mission. Six seems to be a magic number, as that size distributes tasks well and maintains beneficial group cohesion. A number smaller would likely be more burdensome on crew members, possibly affecting the completion of their projects; a number larger would make the Hab seem much more crowded.

Research Report – December 2nd

Crew 216 Research Report 2-Dec-2019
Sol 8
Submitted by: Whittle

1) Psych Scale (TAMU IRB2019-1042D) – 3rd stage surveys to take place on Wednesday.

2) Medical Casts – Nothing further to report.

3) In-Situ Fluorescent Mineral Prospecting – Two EVAs requested for Tues 03-Dec-2019.

4) MDRS Mapping – Nothing further to report (awaiting return of map with desired annotations).

5) MDRS Communications – Summited location on North Ridge, found ideal site for communications rebroadcast point. Conducting GIS analysis of view and tracks, considering maintainability and operational considerations for final report.